40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1199     Entry time: Sun Dec 21 13:00:06 2008
Author: steve 
Type: Update 
Category: all down cond. 
Subject: vac and laser back on 
There was a power outage sometimes early Saturday.

All things are down.

The Maglev was started and reached normal operating condition.
V1 was opened at P1 3.8 mTorr and cc1 is back to 3e-6 Torr now

The MOPA was turned on.
Ion pump HV on to FSS cavity.
ELOG V3.1.3-